OpenStudy (anonymous):

how do you prove that the sum of the squares of two numbers is always greater than or equal to twice their product?

OpenStudy (anonymous):

Could you kindly help me lokisan

OpenStudy (anonymous):

? on your question?

OpenStudy (anonymous):

My problem is posted at http://openstudy.com/groups/mathematics#/updates/4d92ba770b9d8b0bd98b08a9

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

no help?

OpenStudy (anonymous):

Yes I understand taht

OpenStudy (anonymous):

Something strange is happening

OpenStudy (anonymous):

Hello vern, I thought someone was helping you. Consider expanding,\[(a+b)^2\]

OpenStudy (anonymous):

The answers are getting distributed

OpenStudy (anonymous):

Oh..

OpenStudy (anonymous):

Vern, I can help you in a minute.

OpenStudy (anonymous):

@lokisan, you can continue with vern, as my problem is solved

OpenStudy (anonymous):

Thanks a lot, and I must say you are a genius

OpenStudy (anonymous):

Okay, vern, this is what you should do: assume what you have to prove is FALSE. Note, you're trying to eventually show that\[a^2+b^2\ge 2ab\]Let's assume that's false. Then the following would be true:\[a^2+b^2<2ab\]but then, \[a^2+b^2-2ab<0\]that is,\[(a-b)^2<0\] understanding that (a-b)^2 = a^2 + b^2 -2ab. But this conclusion is false, since this is saying that the square of a real number is negative (you can't have such a thing). Therefore, our assumption that \[a^2+b^2<2ab\]leads to a contradiction. We must therefore conclude that the assumption that lead us to this contradiction is false, which means\[a^2+b^2 \ge 2ab\] is true.

OpenStudy (anonymous):

great! thanks